# Mitigation of multi-path propagation artefacts in acoustic targets with cepstral adaptive filtering

### v1

@@ VI Conclusion
The performance of the filtering method was evaluated through two experiments: (i) attenuation of multi-path reflections and motion artefacts in aircraft noise with simulated movement; and, (ii) application to a practical ship-type classification task employing underwater acoustic signals. In the simulated scenario, the filtering strategy effectively recovered the harmonic structure of the target signal across a range of velocities, mitigating both motion artefacts and multi-path effects. In environments affected by the LME, metrics like SNR, IS and LSD improved across all evaluated velocities. In the classification task, the proposed filtering led to improvements in the MCC of 3.46% and 3.29% for the DeepShip and VTUAD V2 datasets, respectively.
Despite the strong results in attenuating motion artefacts, the SNR decreased in cases where no reflected components were present, suggesting that the filtering process may alter signal amplitude. This highlights the need for future improvements in amplitude preservation. Nevertheless, the harmonic structure of the signal was enhanced, as evidenced by a reduction in the IS divergence ranging from 15% to 88% across different velocities.

### v2

@@ VI Conclusion
The performance of the filtering method was evaluated through two experiments: (i) attenuation of multi-path reflections and motion artefacts in aircraft noise with simulated movement; and, (ii) application to a practical ship-type classification task employing underwater acoustic signals. In the simulated scenario, the filtering strategy effectively recovered the harmonic structure of the target signal across a range of velocities, mitigating both motion artefacts and multi-path effects. In environments affected by the LME, metrics like SNR and LSD improved across all evaluated velocities. Also, the improvement over a fixed band filter was between 4-8% across velocities. In the classification task, the proposed filtering led to improvements in the MCC of 3.46% and 3.29% for the DeepShip and VTUAD V2 datasets, respectively.
Despite the strong results in attenuating motion artefacts, the SNR was practically maintained in cases where no reflected components were present, suggesting that the filtering process may alter signal amplitude. This highlights the need for future improvements in amplitude preservation. Nevertheless, the harmonic structure of the signal was enhanced, as evidenced by a reduction in the LSD ranging from 15% to 25% across different velocities. Additionally, the adaptive filtering strategy can represent a computational overhead in comparison with the fixed band filter, needing evaluation on time-restricted applications.